# Selecting Critical Scenarios of DER Adoption in Distribution Grids Using Bayesian Optimization

### v1

@@ Abstract
Index Terms:
Scenarios, PV Adoption, Bayesian Optimization, Gaussian Process
@@ I-A Literature Review
Long-term forecasts of loads and DERs start at the system level to provide a broad picture of load and DER growth across the utility’s territory while ensuring compatibility with other planning processes [6, 3]. Utilities capture data on historical consumption, demographic and economic indicators, weather trends, and DER policy adoption rates [2, 7] and use various statistical-based methodologies to generate such aggregate-level forecasts. For example, [8] introduces a probabilistic forecasting model to predict the magnitude and timing of peak electricity demand, [9] proposes a density forecasting approach to address the uncertainty associated with long-term planning, and [10] uses hourly data to model long-term trends in probabilistic forecasts. A methodology that uses an eXtreme Gradient Boosting algorithm with different sequential configurations is proposed in [11] to combine energy consumption and peak power demand forecasting. In a different direction, the work of [12] proposes a hybrid approach that combines top-down features (economic and weather) with bottom-up features (individual consumer loads and DER adoption propensity) to provide long-term forecasts.
For distribution system planning, utilities disaggregate these system-level forecasts into smaller geographic units using tools like LoadSEER [6, 2]. These tools rely on statistical spatial regressions [13] or agent-based models [14, 3] to generate future scenarios of load growth and DER diffusion. In the literature, earlier works proposed land-use pattern recognition techniques to anticipate new demand using cellular data [15, 16], while recent studies focus on forecasting a combination of loads and DERs, including their temporal characteristics [17, 18]. However, these models are often based on geographical representations (e.g., neighborhoods, zip codes), which do not necessarily correspond to specific feeder locations. The allocation of DER forecasts to individual consumers can be based on rules (e.g., assuming DER power proportional to base load [19]) or on individual consumer models [20]. For example, [21] develops a feeder-specific agent-based model integrating customer GIS data, EV adoption criteria (e.g., neighbor influence, car age), and charging station placement. This feeder-level adoption is highly granular, down to individual parcels and specific customer distribution transformers. In the utility space, Portland General Electric (PGE) developed its own AdopDER tool, a statistical model that captures consumer-level propensity to adopt DERs [22].
@@
Nonetheless, a major limitation of these approaches is that they identify the most critical scenarios in relation to a single equipment (e.g., substations or service transformers). Expanding to the entire distribution system would require identifying all critical scenarios corresponding to all potential system-level violations. In distribution grids with thousands of nodes and millions of potential scenarios of DER adoption, this becomes a search in a high dimensional space. Exhaustively checking scenarios is intractable or at least highly inefficient due to the associated burden of running thousands of power flow evaluations. To resolve this gap, we propose a guided “smart” search via Bayesian Optimization.
@@ I-B Contributions
Statistically, our methodology leverages Gaussian Process (GP) surrogates to efficiently forecast line and bus voltage violations. GPs have been previously used in several power system applications, for instance, to describe the uncertainty of optimal power flows [26], to perform transient stability analysis of large-scale power systems [27, 28], to simulate renewable energy power generation [29, 30], etc. However, this paper is, to the best of our knowledge, the first application of GPs in long-term forecasts of distribution grid load for the purpose of planning.
@@ II-A Scenarios of Adoption
We consider a radial distribution grid with $L$ lines and $B$ buses. Let $\mathcal{A}$ be the set of $A=|\mathcal{A}|$ potential PV adopters, i.e. network nodes where a PV system can be installed. We simulate scenarios of PV adoption with an agent-based extension of the dGen diffusion model [31]. The probability of an agent adopting at time $t+\Delta t$ given the network state at time $t$ is
$\mathbb{P}\!\left(a_{t+\Delta t,j}=1|a_{t,j}=0\right)=\left(p+\dfrac{q}{A}\sum_{n\in\mathcal{A}}a_{t,n}\right),$ (1)
where $p$ and $q$ are the coefficients of innovation and adoption respectively, and $a_{t,j}=1$ if agent $j$ has adopted PV at time $t$, and $a_{t,j}=0$ otherwise. The initial adoption status $a_{0,j}$ is uniformly sampled with a probability matching currently observed adoption rates. For a given time horizon $T$, scenarios describing potential future PV adoption over the network can be obtained by Monte Carlo simulations with transition probabilities (1).
@@ II-B Measuring Distribution Grid Impact
We represent this grid stress as a vector-valued function $\mathbf{f}$ of the scenario $\mathbf{x}$ and evaluate it at both buses and lines as function of potential planning solutions. For buses, we group them into $P$ clusters $\mathcal{B}_{1},\ldots,\mathcal{B}_{P}$ that partition $\mathcal{B}$, representing voltage control zones within the distribution grid. The voltage violation in each cluster (or voltage zone) is measured by the maximum deviation of the computed bus voltage—measured per unit (p.u.) relative to the nominal system voltage—from its acceptable operating range (often ±$\pm 5\%$). This measure quantifies the magnitude of the voltage control problem in each zone, which can be addressed by planning solutions, such as placement of capacitor banks (for undervoltage problems) or voltage regulators (for both over and undervoltage problems). The voltage zones (and number of clusters) can be defined a priori by the utility or automatically obtained by clustering techniques. The left panel of Figure 1 shows an illustrative partition of buses into 12 clusters created with the Louvain community algorithm [32] for a representative feeder (p4rhs8).
@@ III Proposed Methodology
To search for critical scenarios, we propose an algorithm based on BO [33]. The main spirit of BO is to guide the search for critical scenarios using a surrogate-based acquisition function that balances exploration and exploration. Rather than directly maximizing the $f_{k}$’s which requires extensive power flow evaluations, BO constructs a statistical model that uses previously evaluated scenarios to learn the relationship between scenarios and stress objectives. It then carries out a targeted search that only evaluates scenarios that iteratively maximize the acquisition function $\alpha:\mathcal{X}\rightarrow\mathbb{R}$, approximating the probability of a scenario being critical.
@@
Figure 3: Progression of the Pareto frontier projected onto two representative bus objectives. Each point represents the objective values ($f_{3}(\mathbf{x})$, $f_{8}(\mathbf{x})$) of a scenario $\mathbf{x}\in\mathcal{S}_{n^{\star}}$. Evaluated scenarios $\mathbf{x}_{n}$ (and respective Pareto fronts $\mathcal{P}_{n}$) are color-coded in terms of the step $n$ of the BO algorithm.
@@ III-A Gaussian Process Surrogates for Grid Stresses
Related ideas for BO over combinatorial input spaces include choosing an alternative surrogate that can handle high-dim./combinatorial inputs [34, 35], performing the search in a lower-dimensional and/or continuous embedding [36, 37] or modifying the GP kernel to encode the structure of the objective function with respect to the input space [38, 39].
@@ III-B Acquisition Function
The Pareto front we seek is with respect to $V(\boldsymbol{f})$ and we accordingly design an acquisition function $\alpha(\cdot)$ that guides the search towards critical scenarios. The proposed $\mathcal{\alpha}_{ND}(\cdot)$ (where “ND” stands for non-dominated) proxies the probability of $\mathbf{x}$ being a non-dominated (i.e. critical) scenario with respect to violations $V(\boldsymbol{f})$. We evaluate $\mathcal{\alpha}_{ND}(\cdot)$ jointly on $M$ candidate (unevaluated) scenarios $\mathbf{X}^{\prime}=\begin{bmatrix}\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{M}\end{bmatrix}$ by first drawing $N$ samples $\widetilde{\boldsymbol{f}}_{1}(\mathbf{X}^{\prime}),\ldots,\widetilde{\boldsymbol{f}}_{N}(\mathbf{X}^{\prime})$ of $\boldsymbol{f}(\mathbf{X}^{\prime})|\mathcal{D}_{n}$ using the latest GPs. We then compute the set of critical scenarios $\widetilde{\mathcal{P}}_{i}$ in each of these $N$ samples augmented with the evaluated values $f(\mathbf{x}_{i}),i=1,\ldots,n$ of evaluated scenarios, i.e. $\widetilde{\mathcal{P}}_{i}$ is the Pareto front with respect to $V(\widetilde{\boldsymbol{f}}_{i}(\mathbf{X}^{\prime}))\cup V(\boldsymbol{f}(\mathbf{X}_{n}))$. For a scenario $\mathbf{x}^{\prime}\in\mathbf{X}^{\prime}$, we set $\alpha_{ND}(\mathbf{x}^{\prime};\mathcal{D}_{n})$ to the average number of samples in which $\mathbf{x}^{\prime}$ is critical, that is
$\alpha_{ND}(\mathbf{x}^{\prime};\mathcal{D}_{n}):=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\widetilde{\mathcal{P}}_{i}}(\mathbf{x}^{\prime})\approx\mathbb{P}\!\left(\mathbf{x}^{\prime}\in\mathcal{P}(\mathcal{X})|\mathcal{D}_{n}\right).$ (8)
Since $\widetilde{\mathcal{P}}_{i}$ is based jointly on the sampled objectives and the objectives of evaluated scenarios, candidate scenarios only have a positive $\alpha_{ND}$-value if their GP-sampled stresses dominate the stresses of evaluated scenarios. In particular, the maximizer of $\alpha_{ND}$ is the candidate scenario $\mathbf{x}^{\prime}$ that maximizes the likelihood of improving the current Pareto front $\mathcal{P}(\mathbf{X}_{n})$.
The simulation of Pareto fronts $\widetilde{\mathcal{P}}_{i}$ closely follows the approach of [40], although here we map simulated samples to violations. This logic is different from the Expected Hyper-Volume Improvement (EHVI) criterion popular in multi-objective BO [41, 42] that selects the input which leads to the highest average improvement in the volume of the Pareto front. A limitation of EHVI is that its computation scales super-polynomially with the number of objectives and quickly becomes intractable. The construction of $\alpha_{ND}$ is illustrated in Figure 4, which shows simulated Pareto fronts for two bus objectives (for illustration purposes, we do not apply the violation map $V$ that truncates objective values to 0). The initial Pareto front before BO is shown by a dashed black line. The evaluated Pareto front $\mathcal{P}(S_{n})$ is shown by the thick orange line and the global Pareto front $\mathcal{P}(S_{n^{\star}})$ in the space of scenarios, is the thick green line. Each thin gray line represents a simulated Pareto front $\widetilde{\mathcal{P}}_{i}$ based on samples of $f_{1}$ and $f_{8}$ at $\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{M}$. Then, $\alpha(\mathbf{x}^{\prime}_{i};\mathcal{D}_{n})$ is the average number of times the objectives sampled at $\mathbf{x}^{\prime}_{i}$ lie on a simulated staircase.
Note that it is important to sample the candidate scenarios $\mathbf{X}^{\prime}_{n}$ jointly to incorporate the covariance $\Sigma^{(n)}(\mathbf{X}^{\prime}_{n})$ between unobserved scenarios and obtain samples of $f_{k}(\mathbf{x}^{\prime}_{i})$ that are consistent with one another. The specific values of the number of scenarios $M$ and the number of samples $N$ in (8) need to be fine-tuned. Larger $M$ expands exploration but increases the computational overhead as taking GP samples at $M$ scenarios has cubic cost linked to the underlying covariance matrix inversion. Similarly, larger values of $N$ lead to more accurate $\alpha_{ND}$ (and hence more accurate search) with a linearly growing computational burden.
@@
Figure 4: Comparing initial Pareto front $\mathcal{P}(\mathcal{S}_{0})$ to terminal front $\mathcal{P}(\mathcal{S}_{n^{\star}})$ with $n^{\star}=333$ and current Pareto front $\mathcal{P}(\mathcal{S}_{n})$ where $n=55$. We also plot 30 simulated Pareto fronts $\widetilde{\mathcal{P}}_{i}$ obtained by sampling GPs at $M=500$ unevaluated scenarios.
@@ III-C Search space and stopping criterion
Evaluation of $\alpha_{ND}(\cdot)$ on every scenario in the search space $\mathcal{S}_{n}$ becomes expensive when $|\mathcal{S}_{n}|\gg 10^{3}$. Therefore, we further restrict to a set of $M$ candidate scenarios $\mathbf{X}^{\prime}_{n}$, where $M$ is set to the size of the initial search space $\mathcal{S}_{1}$. As $\mathcal{S}_{n}$ grows, we mitigate the risk of promising scenarios being overlooked by sampling them with probability inversely proportional to the frequency with which they have already has been sampled. Thus, scenarios that have never appeared in $\{\mathbf{X}^{\prime}_{k},k\leq n\}$ so far have higher probability of being sampled into $\mathbf{X}^{\prime}_{n+1}$.
@@
$\begin{split}\mathbb{P}\!\left(\exists\ \mathbf{x}\in\mathcal{P}(\mathcal{X})\setminus\mathbf{X}_{n}\right)&\lessapprox\sum_{\mathbf{x}\in\mathcal{S}_{n}\setminus\mathbf{X}_{n}}\mathbb{P}\!\left(\mathbf{x}\in\mathcal{P}(\mathcal{X})\right)\\ &\approx\sum_{\mathbf{x}\in\mathcal{S}_{n}\setminus\mathbf{X}_{n}}\alpha_{ND}(\mathbf{x};\mathcal{D}_{n}).\end{split}$ (9)
@@
$\tau_{n}:=\sum_{\mathbf{x}^{\prime}\in\mathbf{X}^{\prime}}\alpha_{ND}(\mathbf{x}^{\prime};\mathcal{D}_{n}).$ (10)
@@
9:     $\mathbf{x}^{(n)}\leftarrow\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{S}_{n}}\alpha_{ND}(\mathbf{x};\mathcal{D}_{n})\qquad\qquad$ // see (8)
@@
Secondly, fitting GP hyperparameters at every iteration via minimization of the log-likelihood (line 8) unnecessarily increases the computational burden without yielding significant improvements in model performance. GP hyperparameters tend to stabilize after a few iterations and minor changes to them have minimal impact on the BO search. Therefore, we refit the GP hyperparameters (while still updating the posterior means/standard deviations as new evaluations are added) only every $N_{0}$ steps. Finally, to reduce overhead, we select the top $B=4$ scenarios (rather than just the top-1 in line 9) that maximize $\alpha_{ND}$, evaluating this batch in parallel.

### v2

@@ I-A Literature Review
Long-term forecasts of loads and DERs start at the system level to provide a broad picture of load and DER growth across the utility’s territory while ensuring compatibility with other planning processes [6, 3]. Utilities capture data on historical consumption, demographic and economic indicators, weather trends, and DER policy adoption rates [2, 7] and use various statistical-based methodologies to generate such aggregate-level forecasts. For example, [8] introduces a probabilistic forecasting model to predict the magnitude and timing of peak electricity demand, [9] applies spatial forecasting to assess the impact of load growth in power distribution networks, [10] proposes a density forecasting approach to address the uncertainty associated with long-term planning, and [11] uses hourly data to model long-term trends in probabilistic forecasts. A methodology that uses an eXtreme Gradient Boosting algorithm with different sequential configurations is proposed in [12] to combine energy consumption and peak power demand forecasting. In a different direction, the work of [13] proposes a hybrid approach that combines top-down features (economic and weather) with bottom-up features (individual consumer loads and DER adoption propensity) to provide long-term forecasts. More recently, [14] introduce a method based on system dynamics to forecast long-term electricity consumption by dividing the driving factors of energy use into population, per capita GDP, energy intensity, and energy structure. In [15] a hybrid convolutional neural network and long short-term memory method is proposed to forecast monthly energy consumption for three year planning horizons.
For distribution system planning, utilities disaggregate these system-level forecasts into smaller geographic units using tools like LoadSEER [6, 2]. These tools rely on statistical spatial regressions [16] or agent-based models [17, 3] to generate future scenarios of load growth and DER diffusion. In the literature, earlier works proposed land-use pattern recognition techniques to anticipate new demand using cellular data [18, 19], while recent studies focus on forecasting a combination of loads and DERs, including their temporal characteristics [20, 21]. However, these models are often based on geographical representations (e.g., neighborhoods, zip codes), which do not necessarily correspond to specific feeder locations. To forecast household-level consumption, [22] proposes an ensemble of forecasts based on the historical median distribution among similar households. The forecast of DERs at the consumer site can be based on rules (e.g., assuming DER power proportional to base load [23]) or on individual consumer models [24]. For example, [25] develops a feeder-specific agent-based model integrating customer GIS data, EV adoption criteria (e.g., neighbor influence, car age), and charging station placement. This feeder-level adoption is highly granular, down to individual parcels and specific customer distribution transformers. In the utility space, Portland General Electric (PGE) developed its own AdopDER tool, a statistical model that captures consumer-level propensity to adopt DERs [26].
@@
Nonetheless, a major limitation of these approaches is that they identify the most critical scenarios in relation to a single equipment (e.g., substations or service transformers). Expanding to the entire distribution system would require identifying all critical scenarios corresponding to all potential system-level violations. In distribution grids with hundreds of nodes and millions of potential scenarios of DER adoption, this becomes a search in a high-dimensional space making exhaustive evaluation infeasible. We propose a guided “smart” search via Bayesian Optimization, designed to address this challenge by substantially reducing the number of scenarios that must be evaluated, as demonstrated in feeders with several hundred nodes. While we focus on these realistic test cases, the proposed approach offers a scalable foundation that could be extended to larger networks in future work.
@@ I-B Contributions
Statistically, our methodology leverages Gaussian Process (GP) surrogates to efficiently forecast line and bus voltage violations. GPs have been previously used in several power system applications, for instance, to describe the uncertainty of optimal power flows [30], to perform transient stability analysis of large-scale power systems [31, 32], to simulate renewable energy power generation [33, 34], etc. However, this paper is, to the best of our knowledge, the first application of GPs in long-term forecasts of distribution grid load and stress for the purpose of planning.
@@ II-A Scenarios of Adoption
We consider a radial distribution grid with $L$ lines and $B$ buses. Let $\mathcal{A}$ be the set of $A=|\mathcal{A}|$ potential PV adopters, i.e. network nodes where a PV system can be installed. We simulate scenarios of PV adoption with an agent-based extension of the dGen diffusion model [35], introduced in [36]. The probability of an agent adopting at time $t+\Delta t$ given the network state at time $t$ is
$\mathbb{P}\!\left(a_{t+\Delta t,j}=1|a_{t,j}=0\right)=\left(p+\dfrac{q}{A}\sum_{k\in\mathcal{A}}a_{t,k}\right),$ (1)
where $p$ and $q$ are the coefficients of innovation and adoption respectively, and $a_{t,j}=1$ if agent $j$ has adopted PV at time $t$, and $a_{t,j}=0$ otherwise. The initial adoption status $a_{0,j}$ is uniformly sampled with a probability matching currently observed adoption rates. For a given time horizon $T$, scenarios describing potential future PV adoption over the network are obtained by independent Monte Carlo simulations with transition probabilities (1).
@@ II-B Measuring Distribution Grid Impact
We represent this grid stress as a vector-valued function $\boldsymbol{f}$ of the scenario $\mathbf{x}$ and evaluate it at both buses and lines as function of potential planning solutions. For buses, we group them into $P$ clusters $\mathcal{B}_{1},\ldots,\mathcal{B}_{P}$ that partition $\mathcal{B}$, representing voltage control zones within the distribution grid. The voltage violation in each cluster (or voltage zone) is measured by the maximum deviation of the computed bus voltage—measured per unit (p.u.) relative to the nominal system voltage—from its acceptable operating range (often $\pm 5\%$). This measure quantifies the magnitude of the voltage control problem in each zone, which can be addressed by planning solutions, such as placement of capacitor banks (for undervoltage problems) or voltage regulators (for both over and undervoltage problems). The voltage zones (and number of clusters) can be defined a priori by the utility or automatically obtained by clustering techniques. The left panel of Figure 1 shows an illustrative partition of buses into 12 clusters created with the Louvain community algorithm [37] for a representative feeder (p4rhs8).
@@ II-C Defining Critical Scenarios
II-D Use in Distribution Grid Planning
Given the definitions above, the set of critical DER adoption scenarios in the Pareto frontier captures the worst-case violation patterns across all possible adoption decisions. For distribution planning, this small set can replace the high-dimensional space of all adoption realizations. From a utility perspective, this greatly reduces planning complexity regardless of the methodology used. In less sophisticated environments, where planning is still manual, engineers can easily evaluate this small set of critical scenarios. In more advanced utilities, where optimal expansion planning is applied, these scenarios can serve as inputs to more robust or scenario-based optimization.
@@ III Proposed Methodology
To search for critical scenarios, we propose an algorithm based on BO [38]. The main spirit of BO is to guide the search for critical scenarios using a surrogate-based acquisition function that balances exploration and exploitation. Rather than directly maximizing the $f_{k}$’s which requires extensive power flow evaluations, BO constructs a statistical model that uses previously evaluated scenarios to learn the relationship between scenarios and stress objectives. It then carries out a targeted search that only evaluates scenarios that iteratively maximize the acquisition function $\alpha:\mathcal{X}\rightarrow\mathbb{R}$, approximating the probability of a scenario being critical.
@@
Figure 3: Progression of the Pareto frontier projected onto two representative bus objectives. Each point represents the objective values ($f_{3}(\mathbf{x})$, $f_{8}(\mathbf{x})$) of a scenario $\mathbf{x}\in\mathcal{S}_{n^{\star}}$. Evaluated scenarios $\mathbf{x}_{n}$ (and respective Pareto fronts $\mathcal{P}_{n}$) are color-coded in terms of the step $n$ of the BO algorithm. Non-evaluated scenarios are shown in gray.
@@ III-A Gaussian Process Surrogates for Grid Stresses
Related ideas for BO over combinatorial input spaces include choosing an alternative surrogate that can handle high-dim./combinatorial inputs [39], [40], performing the search in a lower-dimensional and/or continuous embedding [41, 42] or modifying the GP kernel to encode the structure of the objective function with respect to the input space [43, 44].
@@ III-B Acquisition Function
In order to guide the search for scenarios on the Pareto front, we construct an acquisition function $\mathcal{\alpha}_{ND}(\cdot)$ (where “ND” stands for non-dominated) that proxies the probability of scenario $\mathbf{x}$ being non-dominated with respect to $V(\boldsymbol{f})$. Thus, the idea is to have $\alpha_{ND}(\mathbf{x};\mathcal{D}_{n})\approx\mathbb{P}\!\left(\mathbf{x}\in\mathcal{P}(\mathcal{X})|\mathcal{D}_{n}\right)$, so that maximizing $\alpha_{ND}$ corresponds to proposing scenarios that are most likely to improve the current Pareto front $\mathcal{P}(\mathbf{X}_{n})$.
Because it is important to incorporate the covariance structure among unobserved scenarios and rank their potential violations consistently, we evaluate $\mathcal{\alpha}_{ND}$ in parallel over $M$ candidate scenarios $\mathbf{X}^{\prime}_{n}=\begin{bmatrix}\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{M}\end{bmatrix}$. To do so, we first draw $N$ samples of $\boldsymbol{f}(\mathbf{X}^{\prime}_{n})|\mathcal{D}_{n}$ using the latest GPs, denoted as $\widetilde{\boldsymbol{f}}_{1}(\mathbf{X}^{\prime}_{n}),\ldots,\widetilde{\boldsymbol{f}}_{N}(\mathbf{X}^{\prime}_{n})$. These capture the distribution of potential violations forecasted for $\mathbf{X}^{\prime}_{n}$. We then compute the subset of critical scenarios $\widetilde{\mathcal{P}}_{i}\subseteq\mathbf{X}^{\prime}_{n}$ in each of these $N$ samples. Then for a candidate $\mathbf{x}_{m}\in\mathbf{X}^{\prime}_{n}$, we set
$\alpha_{ND}(\mathbf{x}_{m};\mathcal{D}_{n}):=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\widetilde{\mathcal{P}}_{i}}(\mathbf{x}_{m}).$ (8)
Thus, $\alpha_{ND}(\mathbf{x}_{m})$ is the frequency that $\mathbf{x}_{m}$ is critical among $\mathbf{X}^{\prime}_{n}\cup\mathbf{X}_{n}$.
The conditional simulation of Pareto fronts closely follows the approach of [45], although here we map simulated samples to violations. This construction is visualized in Figure 4 which shows 30 simulated Pareto fronts $\widetilde{\mathcal{P}}_{i}$ (gray “staircases”) for two bus objectives $f_{1}$ and $f_{8}$. Then, $\alpha(\mathbf{x}_{m};\mathcal{D}_{n})$ is proportional to the number of times the objectives sampled at $\mathbf{x}_{m}$ lie on the staircase. Note that $\mathbf{x}_{m}$ itself is not being plotted, so visually the same scenario would yield $N$ different corners. Although a larger number of samples $N$ in (8) better accounts for the covariance structure $\Sigma^{(n)}(\mathbf{X}^{\prime}_{n})$, leading to more accurate $\alpha_{ND}$ (and hence more accurate search) it comes at the cost of a higher computational burden. Similarly, jointly sampling from a GP at $M$ non-evaluated scenarios $\mathbf{X}^{\prime}_{n}$ to compute $\alpha_{ND}$ and $\tau_{n}$ becomes prohibitive as $M$ grows.
@@
Figure 4: Initial Pareto front $\mathcal{P}(\mathcal{S}_{0})$ (dashed gray), Pareto front $\mathcal{P}(\mathcal{S}_{n})$ at step n = 55 (orange) and true front $\mathcal{P}(\mathcal{S}_{n^{\star}})$ (green) across two representative objectives $f_{3}(\mathbf{x})$, $f_{8}(\mathbf{x})$. We also plot 30 simulated fronts $\widetilde{\mathcal{P}}_{i}$ (solid gray curves) obtained by sampling the terminal GP across $M=500$ non-evaluated scenarios.
@@ III-C Search space and stopping criterion
Evaluation of $\alpha_{ND}(\cdot)$ on every scenario in the search space $\mathcal{S}_{n}$ becomes expensive when $|\mathcal{S}_{n}|\gg 10^{3}$. Therefore, we further restrict to a set of $M$ candidate scenarios $\mathbf{X}^{\prime}_{n}$, where $M$ is set to the size of the initial search space $\mathcal{S}_{1}$. As $\mathcal{S}_{n}$ grows, we mitigate the risk of promising scenarios being overlooked by sampling them with probability inversely proportional to the frequency with which they have already has been sampled. Thus, scenarios that have never appeared in $\{\mathbf{X}^{\prime}_{k},k\leq n\}$ have higher probability of being sampled into $\mathbf{X}^{\prime}_{n+1}$.
@@
$\begin{split}\mathbb{P}\!\left(\exists\ \mathbf{x}\in\mathcal{P}(\mathcal{X})\setminus\mathbf{X}_{n}\right)&\lessapprox\sum_{\mathbf{x}\in\mathcal{S}_{n}\setminus\mathbf{X}_{n}}\mathbb{P}\!\left(\mathbf{x}\in\mathcal{P}(\mathcal{X})\right)\\ &\approx\textstyle\sum_{\mathbf{x}\in\mathcal{S}_{n}\setminus\mathbf{X}_{n}}\alpha_{ND}(\mathbf{x};\mathcal{D}_{n}).\end{split}$ (9)
@@
$\tau_{n}:=\textstyle\sum_{\mathbf{x}^{\prime}\in\mathbf{X}^{\prime}}\alpha_{ND}(\mathbf{x}^{\prime};\mathcal{D}_{n}).$ (10)
@@
9:  $\mathbf{x}^{(n)}\leftarrow\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{S}_{n}}\alpha_{ND}(\mathbf{x};\mathcal{D}_{n})\qquad\quad$ // see (8)
@@
Secondly, fitting GP hyperparameters at every iteration via minimization of the log-likelihood (line 8) unnecessarily increases the computational burden without yielding significant improvements in model performance. GP hyperparameters tend to stabilize after a few iterations and minor changes to them have minimal impact on the BO search. Therefore, we refit the GP hyperparameters (while still updating the posterior means/standard deviations as new evaluations are added) only every $N_{0}$ steps. Finally, to reduce overhead, we select the top $B=4$ scenarios (rather than just the top-1 in line 8) that maximize $\alpha_{ND}$, evaluating this batch in parallel.